# Masked Diffusion Models are Secretly Time-Agnostic Masked Models and Exploit Inaccurate Categorical Sampling

### v3

@@ 1 Introduction
Diffusion models have been extended to discrete data spaces with principled training and sampling (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2023). Compared to ARMs, they predict all the tokens simultaneously and offer a favorable trade-off between generation quality and sampling efficiency. Recently, masked diffusion models (MDMs), the leading variant among discrete diffusion formulations, are emerging as a promising contender of ARMs (Lou et al., 2023). Recent works (Shi et al., 2024; Sahoo et al., 2024) have simplified MDMs to align with the design space of the diffusion models via continuous-time forward processes, training objectives, and sampling procedures, resulting in a unified view and empirical improvements.
However, we argue that the current understanding of MDMs is still quite limited. Positioned at the intersection of diffusion models and masked models, MDMs inherit both the theoretical principles from diffusion models and the simple mechanism from masked models. It remains to be explored how the diffusion theory supports masked models and whether well-established training and sampling recipes of diffusion models can be adapted to enhance MDMs.
In this paper, we dive deeper into the theoretical essence of MDMs. We highlight our key findings: (1) MDMs, in both training and sampling, are essentially time-agnostic masked models (or order-agnostic auto-regressive models), enjoying 20$\times$ faster sampling and diverging from the design choices of diffusion models. This also justifies the theoretical foundation of masked models as they are equivalent and simpler formulations of the more principled MDMs. (2) We challenge previous claims that MDMs can surpass ARMs in text generation by identifying a hidden but critical numerical issue that reduces the token diversity and renders previous evaluations unfair. After fixing it, we find MDMs significantly lagging behind ARMs in generative perplexity.
@@ 3 Revisiting the Training of MDMs
MDMs are defined and trained by the continuous-time forward process (Eqn. (1)), time-dependent network parameterization (Eqn. (4)) and continuous-time ELBO (Eqn. (5)). However, different from continuous-time diffusion models (Song et al., 2021c), the evolution of $\bm{x}_{t}$ is discrete. In the forward process, any token remains the same until it is masked at some time, and then the masked token no longer changes. The evolution trajectories of $(\bm{x}_{t},t)$ are like pairs of “phenotype" and “genotype", where the continuous changes in time $t$ may not be reflected on the observable traits of $\bm{x}_{t}$.
In this section, we aim to disentangle the internal time variable $t$ and the external traits of the masked sequence $\bm{x}_{t}$ at time $t$ in the training of MDMs.
@@ 3.1 Reformulating the ELBO with the Number of Masked Tokens
From Eqn. (8), the time-dependent network $\bm{\mu}_{\theta}(\bm{x},t)$ implicitly parameterizes a time-independent network $\bar{\bm{\mu}}_{\theta}(\bm{x})$ by aggregating the logarithm at the same $\bm{x}$ but different $t$, which can be seen as mixture of experts. The time $t$ is sampled unevenly so that $\alpha_{t}$ follows a Beta distribution $\mathcal{B}(L-n+1,n)$. This distribution has the mode (peak) $\frac{L-n}{L-1}$ and variance $\frac{n(L-n+1)}{(L+1)^{2}(L+2)}\leq\frac{1}{4(L+2)}$. With a large sequence length $L$, the variance is small and the distribution is concentrated around the mode, as illustrated in Figure 2. Moreover, under the best-performing linear schedule $\alpha_{t}=1-t$ in MDMs (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024), we have $t=1-\alpha_{t}$ and the mode of $t$ is $\frac{n-1}{L-1}$, close to the masked ratio $\frac{n}{L}$. Therefore, the time variable $t$ can be seen as a continuous relaxation and smoothing of the masked ratio.
@@
From Eqn. (7), the sequence NELBO can be expressed discretely with the time-agnostic network $\bar{\bm{\mu}}_{\theta}(\bm{x})$. Therefore, Eqn. (7) can serve as a NELBO of masked models in a straightforward approach: uniformly choose the number of masked tokens $n$ from $\{1,\dots,L\}$, uniformly mask $n$ random tokens in $\bm{x}_{0}$ to obtain $\bm{x}_{n}$, and compute the average cross-entropy loss of $\bar{\bm{\mu}}_{\theta}(\bm{x})$ on these $n$ positions. The weighting $\frac{1}{n}$ in this NELBO resembles the likelihood weighting in diffusion models (Song et al., 2021b; Kingma et al., 2021; Lu et al., 2022a; Zheng et al., 2023b), facilitating maximum likelihood training of masked models. Note that early works on order-agnostic auto-regressive models (Uria et al., 2014; Hoogeboom et al., 2021a) already reveal this weighting from a different perspective [^1] The relation between ELBOs of order-agnostic ARMs and MDMs was also mentioned in a recent work (Ou et al., 2024), while they only consider an originally time-agnostic network instead of mixture-of-experts.. While in the context of masked models, there are few discussions on the ELBO. Discussions on related work are placed in Appendix B.
@@ 3.3 Practical Considerations
While there are theoretically equivalent variants for training MDMs (continuous-time/discrete ELBO, time-conditioned/time-independent network), these choices may have practical implications due to differences in network inputs and loss variances. We present some training comparisons and our attempts to improve training (e.g., variance reduction, flow matching) in Appendix I.1. Overall, all options yield similar performance, and the low-discrepancy sampler (Kingma et al., 2021), when applied to time or the number of masked tokens, can significantly reduce the loss variance.
Note that while several works (Lou et al., 2023; Shi et al., 2024) suggest that MDMs are competitive with ARMs in language modeling (beating GPT-2 (Radford et al., 2019) when measured by test/zero-shot perplexity), a more fair comparison (retraining ARMs with the same configurations, Appendix I.1(Sahoo et al., 2024) indicates that MDMs are only advantageous in language understanding tasks (surpassing ARMs and BERT on the GLUE metric (Wang et al., 2018)).
@@ 4 Revisiting the Sampling of MDMs
In the previous section, we demonstrate how the training of MDMs, from both theoretical and empirical perspectives, can be disentangled with the continuous time variable and behave like masked models. In this section, we turn our attention to the sampling of MDMs, which is also performed in continuous time and seems distinct from masked models. We aim to address its current inefficiency problem as well as establish essential insights into its connection with masked models.
@@ 4.1 Inefficiency of Current Sampling
Given the number of sampling steps $N$, the sampling process involves first discretizing the timesteps as $10=t_{0}<t_{1}<\dots<t_{N}=$ 1, and then performing reverse steps $t_{N}\rightarrow t_{N-1}\rightarrow\dots\rightarrow t_{0}$ according to Eqn. (10). Notable characteristics of MDM’s sampling include: (1) Any mask token can only be unmasked once with no further changes. (2) Each sampling step requires a forward pass through the network $\bm{\mu}_{\theta}$ and conducting at most $L$ times of $|\mathcal{X}|$-dimensional categorical sampling, where $L$ is the sequence length and $|\mathcal{X}|$ is the vocabulary size. (3) The number of sampling steps $N$ can be significantly larger than $L$, and a single sampling step may result in no changes to any token in the sequence. (4) As MDMs are trained with the continuous-time ELBO which assumes an infinite number of reverse steps, sampling with finite $N$ theoretically introduces discretization errors. Sampling with a larger $N$ more faithfully reflects the true generation results of MDMs.
@@ 4.3 Parallel Decoding and High-Order Variants
To reduce the approximation error, we follow the recipes of high-order diffusion solvers (Karras et al., 2022; Zhang & Chen, 2022; Lu et al., 2022b; Zheng et al., 2023a) to develop high-order samplers of MDMs. We propose two variants: one based on extrapolating previous network outputs, and the other utilizing a predictor-corrector method to refine the samples. Algorithms of parallel decoding and high-order variants are presented in Appendix G.1.
@@ 5 Are MDMs Better than ARMs? A Critical Fault in Low-Precision Gumbel-Based Categorical Sampling
Before we proceed to verify the effectiveness of our proposed first-hitting sampler, we have to point out a critical fault in MDMs’ original sampling. As suggested by previous works (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024), MDMs seem to surpass ARMs with a sufficient number of sampling steps when measured by the generative perplexity (Gen PPL) [^5] The evaluation metrics used in this paper are introduced in Appendix H.1., as shown in Figure 7a. However, in this section, we identify for the first time a hidden implementation issue existing in previous codebases that makes this observation questionable.
@@ 5.1 Low Token Diversity under Numerous Sampling Steps
Theoretically, the performance of MDMs is more faithfully reflected with an increased number of sampling steps. Empirically, we also observe a reduction in generative perplexity. Notably, an exceptionally low generative perplexity ($<15$) is achieved when the number of sampling steps approaches 50k.
@@ 5.2 Identifying the Numerical Precision Problem
This low generation quality is unexpected, as theory suggests that the approximation error should decrease with more sampling steps. Consequently, samples generated over numerous steps should reflect the true performance of MDMs. We therefore consider this a hidden implementation issue and investigate further to identify the root cause.
@@
Note that auto-regressive LLMs like Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023) use torch.multinomial for categorical sampling, which is also implemented with the Gumbel-max trick in the low-level C++ code of PyTorch. In contrast to the original MDM sampling process, we find the token-by-token decoding process of ARMs and MDMs (by our first-hitting sampler) does not suffer from notable numerical issues under 32-bit precision (Appendix I.2.2).
@@ 5.3 Categorical Sampling with Truncated Gumbel
To the best of our knowledge, this formulation has not been revealed in previous works. Intuitively, with truncated Gumbel, the original class probabilities $\pi_{n}$ are shifted to $\pi_{n}^{\prime}=\pi_{n}\sum_{i=1}^{n}\beta(i)$. This has two main implications: (1) As $\beta(i)\geq 0$ and $\pi_{n}$ are sorted, if $\pi_{n_{1}}>\pi_{n_{2}}$, then the adjusted class probabilities satisfy $\frac{\pi_{n_{1}}^{\prime}}{\pi_{n_{2}}^{\prime}}>\frac{\pi_{n_{1}}}{\pi_{n_{2}}}$. This indicates that relatively larger probabilities are further amplified, creating an effect similar to lowering the temperature. (2) In the sampling step, the probability of unmasking is adjusted based on the network output, resulting in unequal unmasking probabilities at different positions in a sequence. This means some tokens are prioritized to be unmasked, further reducing the randomness and overall entropy.
@@ 6 A Fair Evaluation of MDMs’ Generation
In this section, we will fairly evaluate the generation performance of MDMs and examine the impact of our proposed sampler and the temperature. Our experiments are based on the codebase of MDLM (Sahoo et al., 2024) which is inherited from SEDD (Lou et al., 2023). We fix the categorical sampling to 64-bit floating-point precision so that the numerical truncation is negligible. We directly use pretrained models (AR, SEDD Absorb, MDLM) provided by MDLM, which share the same network architecture and were trained with the same configuration. Additional experiment details are provided in Appendix H.
@@
After the numerical problem is fixed, the entropy returns to a normal level (5.60$\sim$5.70) for all models. Besides, our sampler can be up to 20$\times$ faster than previous sampling strategies of MDMs in terms of the wall-clock time. Despite the notable speedup, the true generative perplexity of MDMs is revealed to be around 100, significantly lagging behind that of counterpart ARMs ($<40$).
@@ 7 Conclusion
In this work, we advance our understanding of masked diffusion models (MDMs) by revealing their theoretical equivalence to masked models and addressing a hidden implementation issue that compromised the fairness of previous evaluations of MDMs’ generative performance. Our investigation challenges earlier claims that MDMs can surpass ARMs in text generation. Despite our negative findings, we acknowledge that our text-based experiments may inherently favor ARMs, as text naturally follows a left-to-right order that ARMs are better suited to model. We believe that MDMs are potentially well-suited for applications where the data’s order-agnostic nature is a key prior.
@@ H.4 Sampling Details
Appendix I Additional Results
I.1 Training Results
I.1.1 Comparison of Training Variants

### v4

@@ 1 Introduction
Diffusion models have been extended to discrete data spaces with principled training and sampling (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2023). Compared to ARMs, they predict all tokens simultaneously and offer a favorable trade-off between generation quality and sampling efficiency. Recently, masked diffusion models (MDMs), the leading variant of discrete diffusion formulations, are emerging as a promising contender of ARMs (Lou et al., 2023). Recent works (Shi et al., 2024; Sahoo et al., 2024) have simplified MDMs to align with the design space of diffusion models via continuous-time forward processes, training objectives, and sampling procedures, resulting in a unified view and empirical improvements. Positioned at the intersection of diffusion models and masked models, MDMs are considered promising as they inherit both the theoretical principles from diffusion models and the simple mechanism from masked models. Moreover, it is believed that MDMs can outperform ARMs in text generation when measured by the common generative perplexity metric (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024).
However, we argue that the current understanding of MDMs is still quite limited in both theoretical and empirical aspects. In this paper [^1] We present our views in a straightforward manner. The earlier version is available in Appendix I., we conduct a thorough and comprehensive investigation and demonstrate that MDMs are essentially a theoretically and empirically equivalent form of typical masked models while being complicated, inefficient, and numerically unstable:
  1. 1.
    The training objective of MDMs is equivalent to that of masked models, differing only by nuanced likelihood-based loss weighting. The introduction of an additional time variable in the loss function provides little benefit in practice. (Section 3)
  2. 2.
    The sampling process of MDMs, being computationally expensive and inefficient, has a theoretically equivalent alternative (our first-hitting sampler) that is up to 20×20$\times$20 × faster and mirrors the random-order, token-by-token decoding process of masked models. (Section 4)
  3. 3.
    The previously reported superiority of MDMs over ARMs on text generation stems from numerical issues during sampling, rather than genuine advantages. (Section 5)
Based on these findings, we argue that the community should reconsider investing efforts in MDMs. That being said, MDMs do provide theoretical insights and supplementary perspectives to masked models, but in practice, the simpler masked models are not only sufficient but also free from inefficiency and numerical instability issues.
@@ 3 Revisiting the Training of MDMs
MDMs are defined and trained by the continuous-time forward process (Eqn. (1)), time-dependent network parameterization (Eqn. (4)) and continuous-time ELBO (Eqn. (5)). However, different from continuous-time diffusion models (Song et al., 2021c), the evolution of $\bm{x}_{t}$ is discrete. The evolution trajectories of $(\bm{x}_{t},t)$ are like pairs of “phenotype" and “genotype", where the continuous changes in time $t$ may not be reflected on the observable traits of $\bm{x}_{t}$. In this section, we aim to disentangle the internal time variable $t$ and the external traits of the masked sequence $\bm{x}_{t}$ in the training of MDMs.
@@ 3.1 Reformulating the ELBO with the Number of Masked Tokens
From Eqn. (8), the time-dependent network $\bm{\mu}_{\theta}(\bm{x},t)$ implicitly parameterizes a time-independent network $\bar{\bm{\mu}}_{\theta}(\bm{x})$ by aggregating the logarithm at the same $\bm{x}$ but different $t$, which can be seen as an ensemble. The time $t$ is sampled unevenly so that $\alpha_{t}$ follows a Beta distribution $\mathcal{B}(L-n+1,n)$. This distribution has the mode (peak) $\frac{L-n}{L-1}$ and variance $\frac{n(L-n+1)}{(L+1)^{2}(L+2)}\leq\frac{1}{4(L+2)}$. With a large sequence length $L$, the variance is small and the distribution is concentrated around the mode, as illustrated in Figure 2. Moreover, under the best-performing linear schedule $\alpha_{t}=1-t$ in MDMs (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024), the mode of $t$ is $\frac{n-1}{L-1}$, close to the masked ratio $\frac{n}{L}$. Therefore, the time variable $t$ can be seen as a continuous relaxation and smoothing of the masked ratio, and we can directly condition the network on the discretely distributed masked ratio instead of the continuous time while yielding similar performance (Appendix J.1).
@@
From Eqn. (7), the sequence NELBO can be expressed discretely with the time-agnostic network $\bar{\bm{\mu}}_{\theta}(\bm{x})$. Therefore, Eqn. (7) can serve as a NELBO of masked models in a straightforward way: uniformly choose the number of masked tokens $n$ from $\{1,\dots,L\}$, uniformly mask $n$ random tokens in $\bm{x}_{0}$ to obtain $\bm{x}_{n}$, and compute the average cross-entropy loss of $\bar{\bm{\mu}}_{\theta}(\bm{x})$ on these $n$ positions. The weighting $\frac{1}{n}$ in this NELBO resembles the likelihood weighting in diffusion models (Song et al., 2021b; Kingma et al., 2021; Lu et al., 2022a; Zheng et al., 2023b), facilitating maximum likelihood training of masked models. Note that early works on order-agnostic auto-regressive models (Uria et al., 2014; Hoogeboom et al., 2021a) already reveal this weighting from a different perspective [^2] The relation between ELBOs of order-agnostic ARMs and MDMs was also mentioned in a recent work (Ou et al., 2024), while they only consider an originally time-agnostic network instead of mixture-of-experts.. While in the context of masked models, there are few discussions on the ELBO. Discussions on related work are placed in Appendix B.
@@ 3.3 Practical Considerations
While there are theoretically equivalent variants for training MDMs (continuous-time/discrete ELBO, time-conditioned/time-independent network), these choices may have practical implications due to differences in network inputs and loss variances. We present some training comparisons and our attempts to improve training (e.g., variance reduction, flow matching) in Appendix J.1. Overall, all options yield similar performance, and the low-discrepancy sampler (Kingma et al., 2021), when applied to time or the number of masked tokens, can significantly reduce the loss variance.
Note that while several works (Lou et al., 2023; Shi et al., 2024) suggest that MDMs are competitive with ARMs in language modeling (beating GPT-2 (Radford et al., 2019) when measured by test/zero-shot perplexity), a more fair comparison (retraining ARMs with the same configurations, Appendix J.1(Sahoo et al., 2024) indicates that MDMs are only advantageous in language understanding tasks (surpassing ARMs and BERT on the GLUE metric (Wang et al., 2018)).
@@ 4 Revisiting the Sampling of MDMs
In the previous section, we demonstrate how the training of MDMs, both theoretically and empirically, can be disentangled with the continuous time variable and behave like masked models. In this section, we turn our attention to the sampling of MDMs, which is also performed in continuous time and seems distinct from masked models. We aim to address its current inefficiency problem as well as establish essential insights into its connection with masked models.
@@ 4.1 Inefficiency of Current Sampling
Given the number of sampling steps $N$, the sampling process involves first discretizing the timesteps as $10=t_{0}<t_{1}<\dots<t_{N}=$ 1, and then performing reverse steps $t_{N}\rightarrow t_{N-1}\rightarrow\dots\rightarrow t_{0}$ according to Eqn. (10). Notable characteristics of MDM’s sampling include: (1) Any mask token can only be unmasked once with no further changes. (2) Each sampling step requires a forward pass through the network $\bm{\mu}_{\theta}$ and conducting at most $L$ times of $|\mathcal{X}|$-dimensional categorical sampling, where $L$ is the sequence length and $|\mathcal{X}|$ is the vocabulary size. (3) The number of sampling steps $N$ can be significantly larger than $L$, and a single sampling step may result in no changes to any token in the sequence. (4) As MDMs are trained with the continuous-time ELBO which assumes an infinite number of reverse steps, it is theoretically rigorous to employ an equivalently large $N$.
@@ 4.3 Parallel Decoding and High-Order Variants
To reduce the approximation error, we follow the recipes of high-order diffusion solvers (Karras et al., 2022; Zhang & Chen, 2022; Lu et al., 2022b; Zheng et al., 2023a) to develop high-order samplers of MDMs. We propose two variants: one based on extrapolating previous network outputs, and the other utilizing a predictor-corrector method to refine the samples (algorithms in Appendix G.1).
@@ 5 Are MDMs Better than ARMs? A Critical Fault in Low-Precision Gumbel-Based Categorical Sampling
Before we proceed to verify the effectiveness of our proposed first-hitting sampler, we have to point out a critical fault in MDMs’ original sampling implementation. As suggested by previous works (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024), MDMs seem to surpass ARMs with a sufficient number of sampling steps when measured by the generative perplexity (Gen PPL) [^6] The evaluation metrics used in this paper are introduced in Appendix H.1., as shown in Figure 7a. However, in this section, we identify for the first time a hidden numerical issue existing in previous codebases that makes this observation questionable.
@@ 5.1 Low Token Diversity under Numerous Sampling Steps
Empirically, a reduction in Gen PPL is observed by increasing the inference budget. In particular, an exceptionally low Gen PPL ($<15$) is achieved when the number of sampling steps approaches 50k.
@@
This low generation quality is unexpected, as theory suggests that increasing sampling steps should yield lower discretization errors and more faithfully reflect the true model performance. We therefore consider this a hidden implementation issue and investigate further to identify the root cause.
@@ 5.2 Identifying the Numerical Precision Problem
Note that auto-regressive LLMs like Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023) use torch.multinomial for categorical sampling, which is also implemented with the Gumbel-max trick in the low-level C++ code of PyTorch. In contrast, we find the token-by-token decoding process of ARMs and MDMs (by our first-hitting sampler) does not suffer from notable numerical issues under 32-bit precision (illustrations and explanations in Appendix J.2.2).
@@ 5.3 Categorical Sampling with Truncated Gumbel
To the best of our knowledge, this formulation has not been revealed in previous works. Intuitively, with truncated Gumbel, the original class probabilities $\pi_{n}$ are shifted to $\pi_{n}^{\prime}=\pi_{n}\sum_{i=1}^{n}\beta(i)$. This has two main implications: (1) As $\beta(i)\geq 0$ and $\pi_{n}$ are sorted, if $\pi_{n_{1}}>\pi_{n_{2}}$, the adjusted class probabilities satisfy $\frac{\pi_{n_{1}}^{\prime}}{\pi_{n_{2}}^{\prime}}>\frac{\pi_{n_{1}}}{\pi_{n_{2}}}$. This indicates that relatively larger probabilities are further amplified, creating an effect similar to lowering the temperature. (2) In the sampling step, the probability of unmasking is adjusted based on the network output, resulting in unequal unmasking probabilities at different positions in a sequence. This implies that some tokens are prioritized to be unmasked, further reducing the randomness and overall entropy.
@@ 6 A Fair Evaluation of MDMs’ Generation
In this section, we will fairly evaluate the generation performance of MDMs and examine the impact of our proposed sampler and the temperature. Our experiments are based on the codebase of MDLM (Sahoo et al., 2024) which is inherited from SEDD (Lou et al., 2023). We fix the categorical sampling to 64-bit floating-point precision so that the numerical truncation is negligible. We directly use pretrained models (AR, SEDD Absorb, MDLM) provided by MDLM, which share the same network architecture and were trained with the same configuration. Additional experiment details are provided in Appendix H. We display some generated text in Appendix J.3.1 to illustrate the token diversity under different sampling strategies.
@@
After the numerical problem is fixed, the entropy returns to a normal level (5.60$\sim$5.70) for all models. Besides, our sampler can be up to 20$\times$ faster than previous sampling strategies of MDMs in terms of the wall-clock time [^8] The efficiency gains (measured by inference wall-clock time) can depend on many factors and may not be as large as 20x in other settings (analyzed in Appendix J.3).. Despite the notable speedup, the true generative perplexity of MDMs is revealed to be around 100, significantly lagging behind that of counterpart ARMs ($<40$).
@@ 7 Conclusion
In this work, we advance our understanding of masked diffusion models (MDMs) by revealing their theoretical equivalence to masked models and uncovering a hidden numerical issue that compromised the fairness of previous evaluations of MDMs’ generative performance. Our findings challenge earlier claims that MDMs can surpass ARMs in text generation. Despite these negative results, we acknowledge that our text-based experiments may inherently favor ARMs, as text naturally follows a left-to-right order that ARMs are better suited to model. Nevertheless, we believe that MDMs may hold potential for applications where an order-agnostic data structure is a key prior, although in practice, simply using masked models may be a better choice.
Limitations and Broader Impact Our investigation yields primarily negative findings. We do not introduce significant enhancements to MDMs, such as network architecture modifications or alternative training strategies. Instead, we negate MDMs on text generation and further question the necessity of MDMs over simpler masked models in other domains. As MDMs are so far the best-performing discrete diffusion models, our study may raise broader concerns regarding the potential of the whole discrete diffusion model family.
@@ H.4 Sampling Details
Appendix I Old Version of the Introduction
We highlight our key findings: (1) MDMs, in both training and sampling, are essentially time-agnostic masked models (or order-agnostic auto-regressive models), enjoying 20$\times$ faster sampling and diverging from the design choices of diffusion models. This also justifies the theoretical foundation of masked models as they are equivalent and simpler formulations of the more principled MDMs. (2) We challenge previous claims that MDMs can surpass ARMs in text generation by identifying a hidden but critical numerical issue that reduces the token diversity and renders previous evaluations unfair. After fixing it, we find MDMs significantly lagging behind ARMs in generative perplexity.
Appendix J Additional Results
J.1 Training Results
J.1.1 Comparison of Training Variants